Working capital: Mukhopadhya Network Associates has a current ratio of 1.60, where the current ratio is defined as follows: current ratio = current assets/current liabilities. The firm’s current assets are equal to $1,233,265, its accounts payables are $419,357, and its notes payables are $351,663. Its inventory is currently at $721,599. The company plans to raise funds in the short-term debt market and invest the entire amount in additional inventory. How much can notes payable increase without the current ratio falling below 1.50?

Please give me the step by step break down of how to get the answer
Answer #1

Let Current Asset be X and Current Liabilities be Y

X/Y = 1.6

Now according to the accounting equation

Asset = Liability + Equity

Now suppose we increase Z amount of inventory by taking short term debt so that our current ratio remains at 1.5

Then the new equation becomes:

X+Z = 1.5(Y+Z)

Here we have considered that equity will remain the same.

Substituting X in the above equation we get Y = 5 Z

Now let us substitute the value of Y from the question which is initial liability i.e 1233265/1.6 = 770790

From the above we get Z = 1,54,158

So we can raise up to 1,54,158 short term debt so that the current ratio does not deep below 1.5.